tommy is twice is old as his younger sister. the sum of the siblings ages is 21. what equation can be used to determine the age of tommys younger sister?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)


tommy is twice is old as his younger sister. the sum of the siblings ages is 21. what equation can be used to determine the age of tommys younger sister?
------
t = 2s
t + s = 21
------
Equation::
(t/2) + s = 21
---------------------
Cheers,
Stan H.
-----------

Answer by ikleyn(52803)

.
x + 2x = 21.